Summary It is cloudy tonight. Totally, 2~hours scientific data taken. J0203+0012 H+K spectra 40min SN 2008hy J+H imaging 34.5min 50.0min J104226 H+K spectra, 15 min (bad condition)

12:38
Nice sunset, a lot of could in the western sky.

00:30 It is totally cloudy.

4:30 We are open. AGW seems dead, restart it.

5:00 AGw is working now, after restart and applying a pointing correction. Even after the correction and only moving a short distance, the guide star is very near the edge of the field.

5:21 begin science observations on J0203. Object is at y~814. First image is #64. Seeing is 0.6'' with some light passing clouds. First image possibly has clouds because object is not visible and sky seems bright. But spectrum is clearly visible in difference of 2nd and 3rd images. Last image is #71. Total of 2400s exposure.

6:10 started lamps. noticed near the end that flexure comp was off, changed it to on and ran again. the correct lamps are #79-#81. #84 is an extra Xe lamp.

6:20 started flats. #85-#90

JMH: The huge coma reported below is because active optics was still running during the lamps above.

6:27 moved to telluric. No guide star found. checking pointing. Off by 20'' in each coordinate. Fixed and back to telluric, found guide star but huge coma. We turned off collimation and the star is in the slit so we're going to take the spectrum anyway. During acquisition the Instrument Control was stuck with the yellow background and the MOS state said moving. There were no error messages, so we arbitrarily moved one of the filter wheels and hit Commit. The move was successful and after it the GUI showed a normal state. So it appears to have just been in the interface.

7:00 Moved to SN2008hy. The script failed as it would have moved the guide probe to an invalid position. We found that the rotation angle should be 90 for the guide star, but it was set to 270. Changed it and now collimating, seeing is 0.5''. First image is #95. Also changed exposures from 30x3 to 15x6 because of bright stars in the field. Seeing is good but still some passing clouds.

7:57 J band finished. Switching to H. First image #118. Still passing clouds so remaining on this imaging program. The seeing is 0.8

9:14 Do pointing. Go to J0958, and we can not find the guide star. Check the pointing and go back, still can not find guide star Go to J1042, and we can find the guide star, and the guider failed in the 4th exposure. We can not see the signal. 10:20 We try to re-acquisition, and failed again. Restart the GCS center, then we are fine

11:05 We go back to J1042, and it is so cloudy and I can only see one star in the field taking 1 minutes exposure.

11:27 We have to close to be safe due to the thick cloud.

12:20 Take darks

-- FuyanBian - 07 Jan 2010
Topic revision: r7 - 07 Jan 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback